# Meta-Homogenization for Knitwear Simulation

### v2

@@ 4. Mesh construction
In mesh construction, we employ the 3D Bresenham algorithm (Žalik et al., 1997; Zhang et al., 2018) to create a volumetric grid, such that each voxel encapsulates a non-zero portion of a yarn thread. To ensure connectivity and eliminate dangling voxels, additional voxels are included in rare cases where the thread passes through a voxel corner. To prevent numerical locking of the hexahedral element (Bathe, 2006), each voxel is further split into six tetrahedrons. Let $n^{Y}$ and $n^{M}$ be the total DOF for the yarns and the mesh, respectively. Typically, $n^{M}$ is around half of $n^{Y}$. In other words, meshing the yarn model is unable to lower the DOF count by orders like in Sperl et al. (2020). As we discuss later, our method excels in a more stable energy formulation and highly parallelizable GPU solver after the volumetric homogenization. This overcomes the cost of high DOF and makes the entire pipeline much more efficient than prior methods (Feng et al., 2024; Sperl et al., 2020).
@@ 5. Shape fitting
With the mesh constructed at the rest pose, our next step is to estimate its shape based on the simulated yarn-level data for each frame. This boils down to finding out the mesh position given the displacement of embedded yarn threads. We use superscripts $(\cdot)^{Y}$ and $(\cdot)^{M}$ to differentiate variables defined on yarns or on the mesh. Let $\bm{x}^{M}\in\mathbb{R}^{3n^{M}}$ be mesh-level position, stacking $x$, $y$, and $z$ coordinates of all the nodes on the tetrahedral mesh. The mesh-to-yarn (M2Y) transfer $\phi^{M2Y}:\mathbb{R}^{3n^{M}}\rightarrow\mathbb{R}^{3n^{Y}}$ is conveniently established with shape-function-based interpolation
(4) $\phi^{M2Y}(\bm{x}^{M})\triangleq\bm{N}\bm{x}^{M}\;,$
where $\bm{N}\in\mathbb{R}^{3n^{M}\times 3n^{Y}}$ contains $N_{i}$ values at yarn endpoints. However, the map along the other direction i.e., from yarn to mesh or Y2M, is less straightforward.
We formulate the Y2M transfer $\phi^{Y2M}$ by minimizing position and deformation inconsistency given an input yarn pose $\bm{x}^{Y}\in\mathbb{R}^{3n^{Y}}$. The deformation error measures the discrepancy of the deformation gradient tensor between yarn and the encapsulating volume mesh. The deformation gradient of a yarn segment is computed as:
@@
(5) $=\bm{R}^{Y}\bm{S}^{Y}=\mathsf{exp}(\bm{\Omega}^{Y})\bm{S}^{Y},$
where $\bar{\bm{x}}_{0,1}^{Y}$ and $\bm{x}_{0,1}^{Y}$ are the rest-shape and deformed positions of two endpoints and $\bar{\bm{n}}_{0,1}^{Y}$ and $\bm{n}_{0,1}^{Y}$ are two mutually perpendicular material normals. Applying the polar decomposition to $\bm{F}^{Y}$ gives a rotation tensor and a symmetric deformation tensor such that $\bm{F}^{Y}=\bm{R}^{Y}\bm{S}^{Y}$, and we use matrix exponential to parameterize the rotation tensor as $\bm{R}^{Y}=\mathsf{exp}(\bm{\Omega}^{Y})$, where $\bm{\Omega}^{Y}$ is a skew-symmetric matrix. Based on this information, we estimate the deformation gradient of the element as
(6) $\bm{F}^{Y2M}(\bm{x}^{Y})=\mathsf{exp}\left(\frac{1}{\sum_{j}L_{j}}\sum_{j}L_{j}\bm{\Omega}_{j}^{Y}\right)\left(\frac{1}{\sum_{j}L_{j}}\sum_{j}L_{j}\bm{S}^{Y}_{j}\right)\;.$
@@
We define the Y2M transfer as the minimizer of the optimization
(7) $\phi^{Y2M}(\bm{x}^{Y})\triangleq\arg\min_{x}\sum_{e}V_{e}\left\|\mathcal{D}\cdot(\bm{B}_{e}\bm{x})-\bm{F}^{Y2M}_{e}\right\|^{2}_{F}\\ +\alpha\left\|\bm{N}\bm{M}^{M}\bm{x}-\bm{M}^{Y}\bm{x}^{Y}\right\|^{2}.$
Here, $\bm{B}_{e}\in\mathbb{R}^{12\times 3n^{Y}}$ is a binary matrix picking 12 DOFs for the tetrahedron $e$ from $\bm{x}$ i.e., $\bm{x}_{e}=\bm{B}_{e}\bm{x}$. $\mathcal{D}\in\mathbb{R}^{3\times 3\times 12}$ is a third-order tensor. It works as a differential operator to compute the deformation gradient of the element out of $\bm{x}_{e}$. $\|\cdot\|_{F}$ denotes the matrix Frobenius norm. $\alpha$ is a hyperparameter, and we set it as $\alpha=0.1$. $\bm{M}^{M}$ is the diagonal lumped mass matrix of the mesh (i.e., using Eq. (2)), and $\bm{M}^{Y}$ is the mass matrix for the yarn. $V_{e}$ is the volume of the element. In other words, Y2M transfer is essentially a volumetric Possion reconstruction (Kazhdan et al., 2006)–we expect $\phi^{Y2M}(\bm{x}^{Y})$ to capture the yarn-level deformation while using $\left\|\bm{N}\bm{M}^{M}\bm{x}-\bm{M}^{Y}\bm{x}^{Y}\right\|^{2}$ as a mass-weighted regularization penalty.
Shape fitting retrieves the information of the mesh from YLS and allows us to estimate inertia forces from an arbitrary yarn motion sequence. As a result, the yarn-level deformation can be understood as a quasi-static one under the non-inertia frame, and the meta-homogenization only needs to focus on fitting the elastic material since mass is decoupled.
@@ 6. Elasticity fitting
Elasticity fitting lies at the core of our meta-homogenization pipeline. In a nutshell, elasticity fitting estimates spatially varying material parameters for each element given input YLS sequences. This is particularly challenging because 1) meta-homogenization seeks a high-dimension material parameter with a large number of knowns, and 2) the optimization takes consideration of all yarn poses. We propose several technical contributions to manage these difficulties, including a second-order Gauss-Newton optimization within the framework of the adjoint method and a harmonic initialization scheme. We borrow the idea of stochastic descent (Bottou et al., 1991) to mitigate the dimensionality concern of space-time optimization.
@@
Elasticity fitting finds the optimal $\bm{\gamma}$ i.e., the value of $\gamma_{e}^{s}$ and $\gamma_{e}^{v}$ for all $n^{E}$ elements, such that M2Y transfer (Eq. 4) of the resultant mesh simulation, $\bm{x}^{M}(\bm{\gamma})$, constitutes a similar yarn dynamics obtained from the full-scale YLS. For a YLS sequence of $n^{F}$ frames
@@
we build an error metric or the loss function between $\bm{x}^{Y}_{i}$ and $\bm{x}^{M}_{i}(\bm{\gamma})$ for $i=1,\cdots,n^{F}$ and minimize the accumulated error for the entire sequence
@@
(11) $\epsilon_{i}(\bm{\gamma})=\sum_{e}V_{e}\left\|\mathcal{D}\cdot\left(\bm{B}_{e}\bm{x}^{M}_{i}(\bm{\gamma})\right)-\bm{F}^{Y2M}_{e}(\bm{x}^{Y}_{i})\right\|^{2}_{F}\\ +\alpha\left\|\bm{N}\bm{M}^{M}\bm{x}^{M}_{i}(\bm{\gamma})-\bm{M}^{Y}\bm{x}^{Y}_{i}\right\|^{2}.$
Here, $\epsilon_{i}$ adopts the same error measure as in Y2M transfer (Eq. 7) but it now depends on $\bm{\gamma}$ since $\bm{x}^{M}_{i}(\bm{\gamma})$ relates to $\bm{\gamma}$ by the mesh-level simulation. The sub-index $i$ here refers to the index of the input $n^{F}$ YLS poses.
Eq. 10 describes a space-time optimization problem. It imposes a significant computational challenge since both $n^{F}$ and $n^{E}$ are big numbers, not to mention $n^{M}$ and $n^{Y}$ are also of high resolution. While first-order descent methods like gradient descent are often preferred due to their simplicity, they fail to deliver a good result even after a large number of iterations in this case. We have to resort to more sophisticated optimization techniques of a higher order for elasticity fitting of Eq. 10.
@@ 6.1. Adjoint Gradient Descent
(12) $\frac{\partial\epsilon_{i}}{\partial\bm{\gamma}}=\frac{\partial\epsilon_{i}}{\partial\bm{x}^{M}_{i}}\cdot\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}.$
The specific form of ${\partial\bm{x}^{M}}/{\partial\bm{\gamma}}$ is up to the mesh-level simulation result, which is often formulated as a variational optimization minimizing the sum of the inertia potential $I$ and the elasticity potential $\sum E_{e}$
(13) $\arg\min_{\bm{x}^{M}_{i}}I+\sum_{e}E_{e},$ where $I$ $=\frac{1}{2\Delta t^{2}}\bm{a}^{M^{\top}}_{i}\bm{M}^{M}\bm{a}^{M}_{i}.$
Shape fitting allows us to approximate $\bm{a}_{i}^{M}$ via
(14) $\bm{a}_{i}^{M}\approx\phi^{Y2M}\left(\bm{x}_{i}^{Y}-2\bm{x}^{Y}_{i-1}+\bm{x}^{Y}_{i-2}\right)-\Delta t^{2}\bm{M}^{M^{-1}}\bm{f}^{M}_{i},$
so that it becomes a known vector and depends on $\bm{x}^{Y}_{i}$, $\bm{x}^{Y}_{i-1}$, $\bm{x}^{Y}_{i-2}$, and the external force $\bm{f}^{M}_{i}$. In other words, the Y2M transfer function $\phi^{Y2M}$ and pre-computed nodal mass (Eq. 2) decouple $I$ from the simulation, converting a dynamic problem to a quasi-static one.
@@
(15) $\bm{g}(\bm{x}^{M}_{i},\bm{\gamma})\triangleq\underbrace{\bm{M}^{M}\phi^{Y2M}\left(\bm{x}_{i}^{M}-2\bm{x}^{Y}_{i-1}+\bm{x}^{Y}_{i-2}\right)-\bm{f}_{i}^{M}}_{\text{inertia force + external force}}+\frac{\partial\sum E_{e}}{\partial\bm{x}_{i}^{M}}=0.$
@@
(16) $\frac{\mathrm{d}\bm{g}}{\mathrm{d}\bm{\gamma}}=\frac{\partial\bm{g}}{\partial\bm{x}^{M}_{i}}\cdot\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}+\frac{\partial\bm{g}}{\partial\bm{\gamma}}=0\Rightarrow\frac{\partial\bm{x}^{M}_{i}}{\partial\bm{\gamma}}=-\left(\frac{\partial\bm{g}}{\partial\bm{x}^{M}_{i}}\right)^{-1}\cdot\frac{\partial\bm{g}}{\partial\bm{\gamma}}.$
To avoid solving the linear system of $\partial\bm{g}/\partial\bm{x}^{M}_{i}$ for $2n^{E}$ times (recalling $\bm{\gamma}$ is a $2n^{E}$-dimension material vector and, therefore, ${\partial\bm{g}/\partial\bm{x}^{M}_{i}\in\mathbb{R}^{3n^{M}\times 2n^{E}}}$), the adjoint method (Tarantola, 2005; Givoli, 2021) leverages an adjoint state vector $\bm{\lambda}$, and evaluates the gradient of the target loss function via
@@
(18) $\left(\frac{\partial\bm{g}}{\partial\bm{x}^{M}_{i}}\right)\bm{\lambda}=\left(\frac{\partial\epsilon_{i}}{\partial\bm{x}^{M}_{i}}\right)^{\top}.$

### v3

@@ 4. Mesh construction
[…]
In mesh construction, we employ the 3D Bresenham algorithm (Žalik et al., 1997; Zhang et al., 2018) to create a volumetric grid, such that each voxel encapsulates a non-zero portion of a yarn thread. To ensure connectivity and eliminate dangling voxels, additional voxels are included in rare cases where the thread passes through a voxel corner. To prevent numerical locking of the hexahedral element (Bathe, 2006), each voxel is further split into six tetrahedrons. Let $n^{Y}$ and $n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ be the total DOF for the yarns and the volume mesh, respectively. Typically, $n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ is around half of $n^{Y}$. In other words, meshing the yarn model is unable to lower the DOF count by orders like in Sperl et al. (2020). As we discuss later, our method excels in a more stable energy formulation and highly parallelizable GPU solver after the volumetric homogenization. This overcomes the cost of high DOF and makes the entire pipeline much more efficient than prior methods (Feng et al., 2024; Sperl et al., 2020).
@@ 5. Shape fitting
With the mesh constructed at the rest pose, our next step is to estimate its shape based on the simulated yarn-level data for each frame. This boils down to finding out the mesh position given the displacement of embedded yarn threads. We use superscripts $(\cdot)^{Y}$ and $(\cdot)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ to differentiate variables defined on yarns or on the volume mesh. Let $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\in\mathbb{R}^{3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}$ be mesh-level position, stacking $x$, $y$, and $z$ coordinates of all the nodes on the tetrahedral mesh. The mesh-to-yarn (V2Y) transfer $\phi^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}2Y}:\mathbb{R}^{3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\rightarrow\mathbb{R}^{3n^{Y}}$ is conveniently established with shape-function-based interpolation
(4) $\phi^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}2Y}(\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}})\triangleq\bm{N}\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\;,$
where $\bm{N}\in\mathbb{R}^{3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\times 3n^{Y}}$ contains $N_{i}$ values at yarn endpoints. However, the map along the other direction i.e., from yarn to mesh or Y2V, is less straightforward.
We formulate the Y2V transfer $\phi^{Y2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ by minimizing position and deformation inconsistency given an input yarn pose $\bm{x}^{Y}\in\mathbb{R}^{3n^{Y}}$. The deformation error measures the discrepancy of the deformation gradient tensor between yarn and the encapsulating volumetric mesh. The deformation gradient of a yarn segment is computed as:
@@
where $\bar{\bm{x}}_{0,1}^{Y}$ and $\bm{x}_{0,1}^{Y}$ are the rest-shape and deformed positions of two endpoints and $\bar{\bm{n}}_{0,1}^{Y}$ and $\bm{n}_{0,1}^{Y}$ are two mutually perpendicular material normals. The deformation gradient includes a rotation and stretch component. While the stretch component can be directly averaged to the corresponding element, the rotation component cannot. To manage the averaging of the rotation component, we apply polar decomposition to $\bm{F}^{Y}$, which yields a rotation tensor and a symmetric deformation tensor such that $\bm{F}^{Y}=\bm{R}^{Y}\bm{S}^{Y}$. We parameterize the rotation tensor using matrix exponential as $\bm{R}^{Y}=\mathsf{exp}(\bm{\Omega}^{Y})$. Note that $\bm{\Omega}^{Y}$ is a skew-symmetric matrix, which can be directly averaged using weighted summation. This geometrically corresponds to averaging the rotation axis and rotation angle(Grassia, 1998). Based on this information, we estimate the deformation gradient of the element as
(5) $\bm{F}^{Y2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}(\bm{x}^{Y})=\mathsf{exp}\left(\frac{1}{\sum_{j}L_{j}}\sum_{j}L_{j}\bm{\Omega}_{j}^{Y}\right)\left(\frac{1}{\sum_{j}L_{j}}\sum_{j}L_{j}\bm{S}^{Y}_{j}\right)\;.$
@@
We define the Y2V transfer as the minimizer of the optimization
(6) $\phi^{Y2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}(\bm{x}^{Y})\triangleq\arg\min_{x}\sum_{e}V_{e}\left\|\mathcal{D}\cdot(\bm{B}_{e}\bm{x})-\bm{F}^{Y2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{e}\right\|^{2}_{F}\\ +\alpha\left\|\bm{N}\bm{M}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\bm{x}-\bm{M}^{Y}\bm{x}^{Y}\right\|^{2}.$
Here, $\bm{B}_{e}\in\mathbb{R}^{12\times 3n^{Y}}$ is a binary matrix picking 12 DOFs for the tetrahedron $e$ from $\bm{x}$ i.e., $\bm{x}_{e}=\bm{B}_{e}\bm{x}$. $\mathcal{D}\in\mathbb{R}^{3\times 3\times 12}$ is a third-order tensor. It works as a differential operator to compute the deformation gradient of the element out of $\bm{x}_{e}$. $\|\cdot\|_{F}$ denotes the matrix Frobenius norm. $\alpha$ is a hyperparameter, and we set it as $\alpha=0.1$. $\bm{M}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ is the diagonal lumped mass matrix of the mesh (i.e., using Eq. (2)), and $\bm{M}^{Y}$ is the mass matrix for the yarn. $V_{e}$ is the volume of the element. In other words, Y2V transfer is essentially a volumetric Possion reconstruction (Kazhdan et al., 2006)–we expect $\phi^{Y2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}(\bm{x}^{Y})$ to capture the yarn-level deformation while using $\left\|\bm{N}\bm{M}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\bm{x}-\bm{M}^{Y}\bm{x}^{Y}\right\|^{2}$ as a mass-weighted regularization penalty.
Shape fitting retrieves the information of the mesh from YLS and allows us to estimate inertia forces as defined in Eq. 14, derived from an arbitrary yarn motion sequence. As a result, the yarn-level deformation can be understood as a quasi-static one under the non-inertia frame, and the volumetric homogenization only needs to focus on fitting the elastic material since mass is decoupled.
@@ 6. Elasticity fitting
Elasticity fitting lies at the core of our volumetric homogenization pipeline. In a nutshell, elasticity fitting estimates spatially varying material parameters for each element given input YLS sequences. This is particularly challenging because 1) volumetric homogenization seeks a high-dimension material parameter with a large number of unknowns, and 2) the optimization takes consideration of all yarn poses. We employ several techniques to address these challenges including the use of second-order Gauss-newton, as described by Zehnder et al. (2021), within the framework of the adjoint method. Additionally, we introduce a novel harmonic initialization scheme for initial value guessing. We borrow the idea of stochastic descent (Bottou et al., 1991) to mitigate the dimensionality concern of space-time optimization.
@@
Elasticity fitting finds the optimal $\bm{\gamma}$ i.e., the value of $\gamma_{e}^{s}$ and $\gamma_{e}^{v}$ for all $n^{E}$ elements, such that V2Y transfer (Eq. 4) of the resultant mesh simulation, $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}(\bm{\gamma})$, constitutes a similar yarn dynamics obtained from the full-scale YLS. For a YLS sequence of $n^{F}$ frames
@@
we build an error metric or the loss function between $\bm{x}^{Y}_{i}$ and $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}(\bm{\gamma})$ for $i=1,\cdots,n^{F}$ and minimize the accumulated error for the entire sequence
@@
(10) $\epsilon_{i}(\bm{\gamma})=\sum_{e}V_{e}\left\|\mathcal{D}\cdot\left(\bm{B}_{e}\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}(\bm{\gamma})\right)-\bm{F}^{Y2{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}_{e}(\bm{x}^{Y}_{i})\right\|^{2}_{F}\\ +\alpha\left\|\bm{N}\bm{M}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}(\bm{\gamma})-\bm{M}^{Y}\bm{x}^{Y}_{i}\right\|^{2}.$
Here, $\epsilon_{i}$ adopts the same error measure as in Y2V transfer (Eq. 6) but it now depends on $\bm{\gamma}$ since $\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}(\bm{\gamma})$ relates to $\bm{\gamma}$ by the mesh-level simulation. The sub-index $i$ here refers to the index of the input $n^{F}$ YLS poses.
Eq. 9 describes a space-time optimization problem. It imposes a significant computational challenge since both $n^{F}$ and $n^{E}$ are big numbers, not to mention $n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ and $n^{Y}$ are also of high resolution. While first-order descent methods like gradient descent are often preferred due to their simplicity, they fail to deliver a good result even after a large number of iterations in this case. We have to resort to more sophisticated optimization techniques of a higher order for elasticity fitting of Eq. 9.
@@ 6.1. Adjoint Gradient Descent
(11) $\frac{\partial\epsilon_{i}}{\partial\bm{\gamma}}=\frac{\partial\epsilon_{i}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\cdot\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}.$
The specific form of ${\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}/{\partial\bm{\gamma}}$ is up to the mesh-level simulation result, which is often formulated as a variational optimization minimizing the sum of the inertial potential $I$ and the elasticity potential $\sum E_{e}$
(12) $\arg\min_{\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}I+\sum_{e}E_{e},$ where $I$ $=\frac{1}{2\Delta t^{2}}\bm{a}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}^{\top}}_{i}\bm{M}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\bm{a}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}.$
Shape fitting allows us to approximate $\bm{a}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}$ via
(13) $\bm{a}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\approx\phi^{Y2{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\left(\bm{x}_{i}^{Y}-2\bm{x}^{Y}_{i-1}+\bm{x}^{Y}_{i-2}\right)-\Delta t^{2}\bm{M}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}^{-1}}\bm{f}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i},$
so that it becomes a known vector and depends on $\bm{x}^{Y}_{i}$, $\bm{x}^{Y}_{i-1}$, $\bm{x}^{Y}_{i-2}$, and the external force $\bm{f}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$. In other words, the Y2V transfer function $\phi^{Y2{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}$ and pre-computed nodal mass (Eq. 2) decouple $I$ from the simulation, converting a dynamic problem to a quasi-static one.
@@
(14) $\bm{g}(\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i},\bm{\gamma})\triangleq\underbrace{\bm{M}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\phi^{Y2{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}\left(\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y}}-2\bm{x}^{Y}_{i-1}+\bm{x}^{Y}_{i-2}\right)-\bm{f}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}_{\text{inertia force + external force}}+\frac{\partial\sum E_{e}}{\partial\bm{x}_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}}=0.$
@@
(15) $\frac{\mathrm{d}\bm{g}}{\mathrm{d}\bm{\gamma}}=\frac{\partial\bm{g}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\cdot\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}+\frac{\partial\bm{g}}{\partial\bm{\gamma}}=0\Rightarrow\frac{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}{\partial\bm{\gamma}}=-\left(\frac{\partial\bm{g}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\right)^{-1}\cdot\frac{\partial\bm{g}}{\partial\bm{\gamma}}.$
To avoid solving the linear system of $\partial\bm{g}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}$ for $2n^{E}$ times (recalling $\bm{\gamma}$ is a $2n^{E}$-dimension material vector and, therefore, ${\partial\bm{g}/\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}\in\mathbb{R}^{3n^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}\times 2n^{E}}}$), the adjoint method (Tarantola, 2005; Givoli, 2021) leverages an adjoint state vector $\bm{\lambda}$, and evaluates the gradient of the target loss function via
@@
(17) $\left(\frac{\partial\bm{g}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\right)\bm{\lambda}=\left(\frac{\partial\epsilon_{i}}{\partial\bm{x}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}}_{i}}\right)^{\top}.$